FOURIER TRANSFORM INFRARED SPECTRA OF FORMALDOXIME: THE ν_{7}, ν_{9}, ν_{11}, ν_{12}, AND 2ν_{12}, BANDS

\underline{G. DUXBURY}^{a}, G. RITCHIE, Department of Physics and Applied Physics, Strathclyde University, Glasgow G4 0NG,Scotland,UK; R. A. BANNAI, Department of Physics, University of Bahrain, PO. Box: 32038, Bahrain; S. KLEE, Physikalisch-Chemisches Institute der Justus-Liebig Universität, D-6300 Giessen, Germany.

A first full rotational analysis of high resolution gas phase spectra of the lowest frequency fundamental bands, ν_{9} and ν_{12}, of formaldoxime, CH_{2}NOH, has been completed. In addition the weak ν_{7} fundamental has been analysed at medium resolution, and the complex structure of the bands in the ν_{11} and 2ν_{12} region has been revealed. It is shown that the complexity of the spectrum of the latter region is due to Coriolis interaction between the overtone 2ν_{12} and the fundamental band ν_{11}. The region covered in this analysis is from 8-30 μm. Since the molecule has Cs symmetry, in principle all the bands of A′ symmetry are AB hybrids. Although the ν_{4}, ν_{5}, ν_{6}, and ν_{8} fundamental bands described previously are predominantly A type, the ν_{7} band is found to be almost a pure B-type band. The ν_{11} and ν_{12} bands are C-type bands . The band origins (cm^{−1}) are ν_{7} 1157.282 ν_{9} 530.01785 ν_{11} 772.78, ν_{12} 397.68499 and 2ν_{12} 778.517.

avisiting fellow at JILA